8 16 14 10 whats the next 3 numbers in the sequence
Answer by HyperBrain(694)

8 16 14 10
the first and third terms are 8 and 14. Note that 14=8+6
the second and fourth terms are 16 and 10. Note that 10=16-6
---
So, the nth term=(n-2)th term + 6 if n is odd
But, if n is even, nth term =(n-2)th term -6.
---
The next term is the 5th term.
Since 5 is odd,
5th term=(5-2)th term + 6=3rd term+6=14+6=20